OpenStudy (qaisersatti):

midterm formula for Q1 please...

OpenStudy (anonymous):

3midterm q1 current through R1 =current through R2 =(V1-R3(I1-I2))/(R1+R2+R3) current through R3 =(V1-(R1R2)(I1-I2))/(R1+R2+R3) 1. Va=(current through R2)*R2+Vb 2. Vb=(current through R3 )*R3 3. power in R1= current through R1*current through R1*R1 4. power in R2=current through R2*current through R2*R2 5. power in R3=current through R3*current through R2*R3 6. power in V1=V1*current through R1 7. power in i1=Vb*current through R3 8. power in i2=-Vb*current through R3

OpenStudy (benzo):

Qaisersatti please me.....The elements in this circuit have the following values: R1=1000Ω, R2=400Ω, R3=600Ω, V1=6V, I1=2A and I2=2A......what is vA, V1, I1, I2

OpenStudy (benzo):

nurali is 32 and -72.......hey help me with please me.....The elements in this circuit have the following values: R1=1000Ω, R2=400Ω, R3=600Ω, V1=6V, I1=2A and I2=2A......what is vA, V1, I1, I2

OpenStudy (benzo):

what the power in V1, I1, I2 formula please